# From Words and Exercises to Wellness: Farsi Chatbot for Self-Attachment Technique

### v2

@@ 3. Proposed Method
Our chatbot follows the same rule-based flowchart as in (Alazraki et al., 2021). For each step in the conversation, we implemented user comprehension modules based on detecting keyword combinations (e.g., to determine negative/positive intent) and classification to identify the user’s current emotion. A detailed illustration of our flowchart is given in Figure 2. Following input comprehension, our framework selects a response from a pool of available utterances based on embedding similarity. The chatbot’s conversation pipeline is shown in Fig. 1.
@@ 3.1. Emotion Classification Module
Fine-grained emotion recognition is a necessary feature for a chatbot offering mental health support. Hence, we implemented an emotion classification module able to identify 12 emotions: Happy, Angry, Anxious, Ashamed, Disappointed, Disgusted, Envious, Guilty, Insecure, Loving, Sad, and Jealous. Due to the scarcity of Farsi datasets labeled for this purpose, we created a new corpus by prompting GPT-3.5 (OpenAI, 2022) in a multi-step fashion to produce diverse Farsi utterances of varying length for each of the 12 emotions. In-context exemplars were deemed unnecessary for this task and thus were not provided to the model. Since Farsi has specific, common expressions to indicate the speaker’s emotion, which were not included in the utterances generated by GPT-3.5, we manually created additional samples, obtaining a final corpus of 6,168 labeled utterances. In Table 1, we show the distribution of the classes in our dataset for emotion classification in Farsi. As evidenced in the table, the classes are approximately balanced. This dataset was used to finetune a ParsBERT model (Farahani et al., 2021) for emotion classification, with results shown in Tab. 2. Our model’s performance (92.13% accuracy, 91.86% F1-score) surpasses previous similar models (Yazdani et al., 2021; Abaskohi1 et al., 2022) while classifying a wider range of emotions, demonstrating significant advancement in the Farsi linguistic domain.
@@ 4.2. User Interface for the Study
Our chatbot was made accessible throughout the evaluation study via a web application optimized for desktop and mobile devices running various operating systems. To further maximize accessibility, the chatbot allowed for both voice and text input and was able to give spoken and written responses. We aimed to make the application suitable for a wide variety of users and ages by allowing the choice of either ‘formal’ or ‘informal’ conversation, leveraging the different pools of utterances that we created for this purpose. In addition, to facilitate usability, we included in the web application the following: (1) information about the study and introductory information about SAT, visible at login; (2) a page for frequently asked questions (FAQ) with in-depth explanations of the core concepts of SAT; (3) the SAT Teacher extension, described in Section 3.4, for answering any further questions about SAT and the chatbot itself on demand. Our analysis shows that over 90% of participants used the SAT Teacher. The platform’s interface is shown in Fig. 3.
@@ 4.4. Results
In line with previous research (Alazraki et al., 2021; Law et al., 2022), we asked the volunteers to evaluate the chatbot’s empathy, engagingness, ability to interpret human emotions correctly, and overall performance. Additionally, we asked them to score the effectiveness of the chatbot at increasing positive emotions, decreasing negative ones, and helping users feel better, as well as the effect of SAT on overall well-being. Lastly, we evaluated users’ satisfaction with the answers provided by the SAT teacher. For all of the above metrics, study participants were asked to assign a discrete score between 1 and 10, where 10 indicated strongest performance. The results of this evaluation are shown in Tab. 3, and a comprehensive breakdown of the scores assigned by the volunteers for each question is given in Figure 4.
In the interest of carrying out a thorough qualitative evaluation as well as a quantitative one, we also collected open-ended user feedback on the chatbot’s main benefits and limitations. We then clustered these comments in a thematic analysis into 13 overarching themes, shown in Tab. 4 and accompanied by a representative feedback sample that we manually selected.
@@ 5.2. Findings
The quantitative feedback collected during the study indicates that the majority of volunteers scored the chatbot positively, with very similar average scores for all evaluation dimensions, as evidenced in Tab. 3. Among the best scoring features were the chatbot’s engagingness (mean = 7.58, STD = 2.29), the overall performance (mean = 7.48, STD = 2.40), and the SAT Teacher’s responses (mean = 7.42, STD = 2.29). The accuracy of the emotion recognition module, while receiving slightly lower scores than other metrics (mean = 6.78, STD = 2.84), was still deemed adequate by most participants. While largely positive, these scores, ranging between 6.78 and 7.58 on a scale from 1 to 10, show that the platform could be further improved. Indeed, we found that while the majority of study participants rated the platform positively (i.e., with a score between 6 and 10), approximately a quarter assigned a low score (i.e., a score between 1 and 5) to each point in the evaluation questionnaire, as evidenced in Fig. 4.
Among the qualitative feedback (Tab. 4), the helpfulness of the chatbot was most often mentioned (16 comments), followed by its user-friendliness (10 comments) and the beneficial effects of the recommended SAT exercises (nine comments). While negative comments were overall less frequent than positive ones, several users disliked the semi-structured nature of the conversation and stated they would prefer to be able to talk to the chatbot in a completely open-ended manner (eight comments). Other, less frequent themes that may have negatively impacted the perception of our platform revolved around (1) the difficulty of some of the SAT exercises (noted by five users), (2) the rigidity of the chatbot’s responses (noted by four users); (3) insufficient information and guidance to successfully complete the SAT exercises (noted by four users); (4) the perceived slowness of the platform (noted by three users); (5) lack of images that could otherwise make the material more engaging, particularly in exercise descriptions (noted by three users); (6) not enough empathy toward the user throughout the conversation (noted by three users).

### v3

@@ 3. Proposed Method
Our chatbot follows the same rule-based flowchart as in (Alazraki et al., 2021). For each step in the conversation, we implemented user comprehension modules based on detecting keyword combinations (e.g., to determine negative/positive intent) and classification to identify the user’s current emotion. A detailed illustration of our flowchart is given in Figure 2. Following input comprehension, our framework selects a response from a pool of available utterances based on embedding similarity. The chatbot’s conversation pipeline is shown in Figure 1.
@@ 3.1. Emotion Classification Module
Fine-grained emotion recognition is a necessary feature for a chatbot offering mental health support. Hence, we implemented an emotion classification module able to identify 12 emotions: Happy, Angry, Anxious, Ashamed, Disappointed, Disgusted, Envious, Guilty, Insecure, Loving, Sad, and Jealous. Due to the scarcity of Farsi datasets labeled for this purpose, we created a new corpus by prompting GPT-3.5 (OpenAI, 2022) in a multi-step fashion to produce diverse Farsi utterances of varying length for each of the 12 emotions. In-context exemplars were deemed unnecessary for this task and thus were not provided to the model. Since Farsi has specific, common expressions to indicate the speaker’s emotion, which were not included in the utterances generated by GPT-3.5, we manually created additional samples, obtaining a final corpus of 6,168 labeled utterances. In Table 1, we show the distribution of the classes in our dataset for emotion classification in Farsi. As evidenced in the table, the classes are approximately balanced. This dataset was used to finetune a ParsBERT model (Farahani et al., 2021) for emotion classification, with results shown in Table 2. Our model’s performance (92.13% accuracy, 91.86% F1-score) surpasses previous similar models (Yazdani et al., 2021; Abaskohi1 et al., 2022) while classifying a wider range of emotions, demonstrating significant advancement in the Farsi linguistic domain.
@@ 4.2. User Interface for the Study
Our chatbot was made accessible throughout the evaluation study via a web application optimized for desktop and mobile devices running various operating systems. To further maximize accessibility, the chatbot allowed for both voice and text input and was able to give spoken and written responses. We aimed to make the application suitable for a wide variety of users and ages by allowing the choice of either ‘formal’ or ‘informal’ conversation, leveraging the different pools of utterances that we created for this purpose. In addition, to facilitate usability, we included in the web application the following: (1) information about the study and introductory information about SAT, visible at login; (2) a page for frequently asked questions (FAQ) with in-depth explanations of the core concepts of SAT; (3) the SAT Teacher extension, described in Section 3.4, for answering any further questions about SAT and the chatbot itself on demand. Our analysis shows that over 90% of participants used the SAT Teacher. The platform’s interface is shown in Figure 3.
@@ 4.4. Results
In line with previous research (Alazraki et al., 2021; Law et al., 2022), we asked the volunteers to evaluate the chatbot’s empathy, engagingness, ability to interpret human emotions correctly, and overall performance. Additionally, we asked them to score the effectiveness of the chatbot at increasing positive emotions, decreasing negative ones, and helping users feel better, as well as the effect of SAT on overall well-being. Lastly, we evaluated users’ satisfaction with the answers provided by the SAT teacher. For all of the above metrics, study participants were asked to assign a discrete score between 1 and 10, where 10 indicated strongest performance. The results of this evaluation are shown in Table 3, and a comprehensive breakdown of the scores assigned by the volunteers for each question is given in Figure 4.
In the interest of carrying out a thorough qualitative evaluation as well as a quantitative one, we also collected open-ended user feedback on the chatbot’s main benefits and limitations. We then clustered these comments in a thematic analysis into 13 overarching themes, shown in Table 4 and accompanied by a representative feedback sample that we manually selected.
@@ 5.2. Findings
The quantitative feedback collected during the study indicates that the majority of volunteers scored the chatbot positively, with very similar average scores for all evaluation dimensions, as evidenced in Table 3. Among the best scoring features were the chatbot’s engagingness (mean = 7.58, STD = 2.29), the overall performance (mean = 7.48, STD = 2.40), and the SAT Teacher’s responses (mean = 7.42, STD = 2.29). The accuracy of the emotion recognition module, while receiving slightly lower scores than other metrics (mean = 6.78, STD = 2.84), was still deemed adequate by most participants. While largely positive, these scores, ranging between 6.78 and 7.58 on a scale from 1 to 10, show that the platform could be further improved. Indeed, we found that while the majority of study participants rated the platform positively (i.e., with a score between 6 and 10), approximately a quarter assigned a low score (i.e., a score between 1 and 5) to each point in the evaluation questionnaire, as evidenced in Figure 4.
Among the qualitative feedback (Table 4), the helpfulness of the chatbot was most often mentioned (16 comments), followed by its user-friendliness (10 comments) and the beneficial effects of the recommended SAT exercises (nine comments). While negative comments were overall less frequent than positive ones, several users disliked the semi-structured nature of the conversation and stated they would prefer to be able to talk to the chatbot in a completely open-ended manner (eight comments). Other, less frequent themes that may have negatively impacted the perception of our platform revolved around (1) the difficulty of some of the SAT exercises (noted by five users), (2) the rigidity of the chatbot’s responses (noted by four users); (3) insufficient information and guidance to successfully complete the SAT exercises (noted by four users); (4) the perceived slowness of the platform (noted by three users); (5) lack of images that could otherwise make the material more engaging, particularly in exercise descriptions (noted by three users); (6) not enough empathy toward the user throughout the conversation (noted by three users).